OpenStudy (anonymous):

Can anyone help me with this question, at least show me how to set up which equations to solve it? "A shipping company designs a container that is a rectangular prism with a length that is 3 inches more than its width and a height that is 1 inch less than its width. If the volume of the container is 160 cubic inches, find its dimensions. Show all your work."

OpenStudy (phi):

Did you draw a picture and label it ?

OpenStudy (anonymous):

 This is basically what I have drawn, I'm just not completely sure how to go about solving it.

OpenStudy (phi):

the volume is length*width*height or in this case x(x-1)(x+3) = 160 multiply out. what do you get ?

OpenStudy (anonymous):

I really don't want to sound stupid, but I kind of need a refresher on how to do that, like what steps to take to multiply it out and find the dimensions, like solving for x or something.

OpenStudy (phi):

This gives the basics http://www.khanacademy.org/math/algebra/polynomials/multiplying_polynomials/v/multiplication-of-polynomials in this case, first do (x-1)*(x+3) if we do FOIL (first terms), outer terms, inner terms, last terms x*x +x*3 + -1*x + -1*3 or x^2 +2x -3 (the other way is "distribute" the (x-1): (x-1)*x + (x-1)*3 now distribute again: x^2 -x + 3x -3 and combine the x terms: x^2 + 2x-3 now multiply by x (because we are doing x*(x-1)*(x+3) ) we get x*(x^2 + 2x-3) = x^3 +2x^2-3x that is the volume and it equals 160: x^3 +2x^2-3x= 160 or, subtracting 160 from both sides, to put this cubic equation in standard form: x^3 + 2x^2 -3x -160 =0

OpenStudy (phi):

cubics are complicated to solve. see http://en.wikipedia.org/wiki/Cubic_function for the hairy details. I imagine they want you to use a graphing calculator to solve this? or, use wolfram (which may or may not be cheating, depending on what you are learning...) http://www.wolframalpha.com/input/?i=x%5E3+%2B+2x%5E2+-3x+-160+%3D0

OpenStudy (phi):

wolfram found x=5. If we use that in your picture, we find width=5, height = 4 and length= 8

OpenStudy (phi):

as a check: 5*4*8= 20*8= 160 cu. inches which matches what they told us is the volume.

OpenStudy (anonymous):

Thank you so much! That helped tremendously!